The lengths of the sides of a triangle are 3x, 5x – 12, and x + 20. Find the value of x so that the triangle is isosceles.
Answer by mananth(12270)

Any two side are equal in isoscles triangle

5x-12=x+20
4x=32
x=8
5x-12=28
x+20=28
5x=40
28,28,40